OpenStudy (anonymous):

how would I come up with 66 2/3% of 39?

OpenStudy (jdoe0001):

\(\bf {\color{red}{ 66\frac{2}{3}}}\%\ of \ 39=\cfrac{{\color{red}{ 66\frac{2}{3}}}}{100}\times 39\)

OpenStudy (texaschic101):

66 2/3 % = 66.67% 66.67% = .6667 " of " means multiply .6667 (39) = 26.0013